OpenStudy (anonymous):

Use the associative property to finish the equation: (4x + 2) + x = A. 4x + (2 + x) B. (2 + 4x) + x C. x + (4x + 2) D. 4x + (x + 2)

OpenStudy (anonymous):

The right answer is D. For more info on associative property: http://en.wikipedia.org/wiki/Associative_property

Directrix (directrix):

A) (4x + 2) + x = 4x + (2 + x) The Associative Property is the grouping property. The order of the variables being grouped does not change. Hence, Answer A is correct. Answer D is an illustration of the Commutative Property. Examples of Associative Property Addition (2 + 3) + 5 = 2 + (3 + 5) http://www.northstarmath.com/sitemap/associativeproperty.html